Let A={a, b, c}, B={d, e}, C={a, d}.

Find (i) A × B (ii) B × A (iii) A × (B C)

(iv) (A C) x B (v) A ∆ (B – C) (vi) Complement of (A ∆ B)

(vii) P(B C)


1
Expert's answer
2021-11-12T15:26:20-0500

i)

"A × B=\{(a,d),(a,e),(b,d),(b,e),(c,d),(c,e)\}"


ii)

"B × A=\{(d,a),(d,b),(d,c),(e,a),(e,b),(e,c)\}"


iii)

"(B C)=\{d\}"

"A × (B C)=\{(a,d),(b,d),(c,d)\}"


iv)

"(A C)=\{a\}"

"(A C) \times B=\{(a,d),(a,e)\}"


v)

"B – C=\{e\}"

"A ∆ (B – C)=(A-(B-C))\cup ((B-C)-A)=\{a, b, c\}\cup \{e\}=\{a, b, c,e\}"


vi)

"A ∆ B=(A-B)\cup (B-A)=\{a, b, c\}\cup \{d, e\}=\{a, b, c,d,e\}"

if universal set:

"U=\{a, b, c,d,e\}"

then complement of (A ∆ B) is empty set "\empty"


vii)

"(B C)=\{d\}"

"P(B C)" is the set of all subsets of (BC)

"P(B C)=\{\{\},\{d\}\}"